# HD-GEN: A High-Performance Software System for Human Mobility Data Generation Based on Patterns of Life

### v1

@@ Abstract.
Understanding individual-level human mobility is critical for a wide range of applications. Real-world trajectory datasets provide valuable insights into actual movement behaviors but are often constrained by data sparsity and participant bias. Synthetic data, by contrast, offer scalability and flexibility but frequently lack realism. To address this gap, we introduce a comprehensive software pipeline for calibrating, generating, processing, and visualizing large-scale individual-level human mobility datasets that combine the realism of empirical data with the control and extensibility of Patterns-of-Life simulations. Our system consists of four integrated components. (1) a data generation engine constructs geographically grounded simulations using OpenStreetMap data to produce diverse mobility logs. (2) a genetic algorithm–based calibration module fine-tunes simulation parameters to align with real-world mobility characteristics, such as daily trip counts and radius of gyration, enabling realistic behavioral modeling. (3) a data processing suite transforms raw simulation logs into structured formats suitable for downstream applications, including model training and benchmarking. (4) a visualization module extracts key mobility patterns and insights from the processed datasets and presents them through intuitive visual analytics for improved interpretability.
GeoLife, Patterns of Life, Simulation, Realistic Trajectory Datasets
$c$s: Information systems Location based servicesccs: Software and its engineering Software design engineeringccs: Computing methodologies Simulation tools
@@ 1. Introduction
Human mobility datasets are fundamental for understanding human behavior (toch2019analyzing; zhu2024generic), mobility patterns (mokbel2018mobility), and traffic dynamics (chen2022lane). They enable a wide range of applications, including outlier and anomaly detection (zhang2024large; zhang2024transferable; amiri2024urban), infectious disease modeling (kohn2023epipol), and urban planning and infrastructure design (isaacman2012human). Mobility datasets can be broadly categorized as real or synthetic. Real world data provide high fidelity but raise serious quality and privacy concerns, while synthetic data offer improved scalability and privacy preservation at the cost of reduced realism. In practice, privacy regulations (xiao2015protecting; khoshgozaran2011location; krumm2009survey), de-identification techniques, and voluntary data collection introduce sparsity and information loss, resulting in significant trade-offs between data utility, realism, and accessibility (zhu2024synmob; amich2025exploring).
@@
To mitigate these limitations, researchers increasingly rely on simulated trajectory datasets that provide scalable, noise free mobility information without privacy risks. However, many existing trajectory simulators rely on random or simplistic destination selection (duntgen2009berlinmod; brinkhoff2002framework) or parametric trip distributions (levandoski2012lars; armenatzoglou2013general). While such approaches are suitable for benchmarking storage systems or query processing, they fail to capture the intentional and routine driven nature of real human mobility.
@@
The Pattern of Life (POL) simulation framework (kim2020location; zufle2023urban; amiri2024patterns) addresses this gap by modeling human needs, activities, and daily routines. POL generates purposeful mobility behaviors such as commuting, dining, and social interactions, producing trajectories that better reflect real world movement semantics. POL has been used to generate large scale datasets with realistic mobility patterns, social interactions, and check-in behavior. Prior work includes the generation of fine grained mobility and social network datasets across multiple U.S. cities (amiri2023massive), datasets with injected anomalies for mobility anomaly detection (amiri2024urban), and a GeoLife calibrated dataset optimized via genetic algorithms to balance realism and scalability (amiri2024geolife+).
@@
In this paper, we extend our short SIGSPATIAL 2025 version (amiri2025hd) and consolidate our prior contributions (amiri2024patterns; amiri2024geolife+; amiri2023massive) into a complete and systematic POL based data generation system. We introduce a unified software platform that integrates trajectory generation, calibration, processing, and visualization into a coherent end to end pipeline, referred to as HD-GEN. The platform expands the functionality of POL while emphasizing reproducibility, scalability, and usability for large scale experimental and operational settings.
@@
  • delivers an improved open source repository ({https://github.com/onspatial/hd-gen-large-scale-human-mobility-generator}) with clearer documentation and integrated visualization toolkits
The remainder of the paper is organized as follows. In Section 2, we review relevant background information and related prior work. In Section 3, we provide a technical overview of POL and describe the software design and system-level details of HD-GEN. In Section 4, we summarize the datasets produced using HD-GEN. Finally, in Section 5, we conclude by summarizing the contributions of this work and discussing directions for future research.
2. Background Information and Related Works
In this section, we provide background information and related works on human mobility data generation, including an overview of the Patterns-of-Life (POL) simulation and a discussion of existing trajectory datasets. We first describe the POL framework, its behavioral foundations, and its suitability for generating large-scale, realistic mobility data. We then briefly review commonly used real and synthetic trajectory datasets and highlight their limitations, motivating the use of POL and the need for calibration against observed mobility patterns.
Pol details
@@ 2.1. The Patterns-of-Life Simulation (POL)
The Patterns-of-Life (POL) simulation, introduced in Urban Life (zufle2023urban), is a city-scale, agent-based framework designed to generate realistic human mobility trajectories and interaction patterns. The simulation instantiates autonomous agents within an urban environment constructed from OpenStreetMap data, including road networks, residential zones, workplaces, and recreational venues. Agents perceive and interact with this environment while independently making decisions that reflect the routines and choices of individuals in everyday life.
Agent behavior in POL is governed by a combination of physiological, financial, and social needs inspired by Maslow’s hierarchy (maslow1943theory), together with goal-oriented decision making based on the Theory of Planned Behavior (ajzen1991theory). At each simulation time step, agents evaluate competing needs such as earning income, acquiring food, maintaining social relationships, and participating in leisure activities. These needs are balanced against practical constraints including time availability, spatial accessibility, and personal resources. The resulting decisions determine when and where agents travel, producing realistic daily schedules and mobility patterns. As agents move between different types of locations, the simulation generates fine-grained spatiotemporal trajectories as well as implicit social graphs that emerge from repeated co-location and interaction. This design makes POL well suited for controlled experimentation and downstream analysis in mobility modeling, social network analysis, and urban studies.
@@
Fig. 1 illustrates the interaction between a person agent and the simulation world. Each person is modeled as an autonomous agent with multiple needs, including food, sleep, shelter, financial safety, and love need from social connections. These needs evolve continuously over time and drive the agent’s decision making process. To satisfy its needs, the agent interacts with the environment, which defines mobility constraints, a spatial network, and points of interest such as workplaces, homes, and public venues. The environment determines where the agent can move and which actions are feasible at each location. The agent perceives the environment, selects actions based on its current needs and context, and executes those actions through movement and interactions with points of interest, such as working at a workplace, or with other agents at recreational sites. All agent states, actions, and environment responses are recorded by the log system. The generated logs capture realistic spatiotemporal behavior and serve as the output of the simulation for downstream analysis and modeling.
@@
The POL simulation is implemented in Java, configurable through a command-line interface, and publicly available on GitHub (zufle2023urban). Users can specify city regions, population sizes, and behavioral parameters, enabling reproducible and extensible experiments. Fig. 2 presents the end to end simulation workflow. The process begins with a parameter sweep, where multiple configurations are defined to explore variations in agent behavior and environmental conditions. These configurations are formalized during parameterization, producing a consistent set of parameters that govern the simulation. In parallel, the map generator produces spatial representations of the environment, including networks and points of interest. The generated maps, together with the selected parameters, are passed to the initialization stage, where agents, environments, and initial states are instantiated. The simulation stage executes agent interactions over time based on the initialized state and active parameters. During execution, manipulation modules can dynamically modify parameters or system states, enabling interventions, scenario changes, or policy testing. The simulation state can be serialized to disk to implement checkpointing, allowing execution to be paused, restored, and reproduced in long running experiments. All simulation outputs are stored in structured formats guided by a predefined schema definition, which specifies how agent states, actions, and environmental events are recorded. The log system collects these records, producing standardized logs that serve as the final output of the simulation.
@@ 2.2. Related Work and Relation to HD-GEN
Human mobility data can be collected from real-world observations or synthesized using computational models. Real-world trajectory datasets capture observed mobility behavior and provide valuable insight into human movement and activity patterns. A widely used example is the GeoLife dataset (zheng2011geolife), which contains four years of high-resolution GPS traces from approximately 180 users in Beijing, spanning both commuting and leisure activities. More recently, the large-scale YJMOB100K dataset (yabe2024yjmob100k) provides anonymized mobility records from 100,000 users at 30-minute intervals, aggregated into 500-meter spatial cells. While this scale enables population-level analysis, the spatial aggregation limits fine-grained, place-based behavioral inference. Additional trajectory datasets focus on vehicles, including taxis (yuan2010t; PSG09) and buses (DC18). These datasets are valuable for modeling transportation systems but do not accurately represent individual human activity patterns, since each trip may involve different occupants and does not reflect continuous personal behavior. Synthetic trajectory datasets address limitations of real mobility data, including privacy constraints, proprietary access, and restricted scale. Recent deep generative approaches learn mobility patterns to produce realistic spatiotemporal trajectories, including factorized models such as EETG (zhang2022factorized) and diffusion-based methods (zhu2024synmob). Optimization-based techniques further aim to enhance incomplete or sparse trajectory data (zhu2024generic; hu2022processing). In contrast, the POL simulation framework (kim2020location; zufle2023urban; amiri2024patterns) generates trajectories driven by modeled human needs and decision processes rather than purely geometric or kinematic realism.
In our prior work (amiri2023massive), POL was used to generate a large-scale dataset exceeding 1.5 TB of mobility, check-in, and social interaction data. We further demonstrated the resulting data generator and its ability to reproduce key mobility and activity patterns in (amiri2024patterns). Although the default parameter settings yield qualitatively realistic behavior, regional variation was not systematically characterized. This limitation motivated our subsequent study (amiri2024geolife+), in which we calibrated POL to better match mobility patterns observed in the GeoLife dataset for Beijing. Finally, we integrated these efforts into a unified framework, HD-GEN, which was introduced in a concise system paper (amiri2025hd).
@@ 3.2. POL Details
This section describes the new features added to POL as part of this extension, including checkpointing and log streaming.
@@
The Pattern of Life simulation is deterministic, meaning that executing the simulation with the same input parameters always produces identical results. The simulation is also portable across different machines and operating systems, allowing execution to continue without breaking the pipeline or altering outcomes. These properties are essential for exploring multiple scenarios while preserving previously executed configurations.
@@ 3.3. HD-GEN Details
To generate simulated data, we developed tools for regional map preparation, simulation execution, and large scale dataset generation in parallel. Map generation relies on OpenStreetMap (OSM) data accessed through the Overpass API to extract building footprints, road networks, and land use information. These raw features are further refined using predictive models to classify building usage. The map generation process produces three shapefiles, buildings.shp, buildingUnits.shp, and walkways.shp, which collectively define the spatial environment used by the POL simulation. Figure 6 shows an example script that generates a new map for the Minneapolis region. In addition to bounding box based queries, the tools support custom Overpass queries, allowing flexible selection of geographic regions and features for map generation.
@@
For calibration, we used the GeoLife dataset as a real-world reference and applied a genetic algorithm guided by a similarity function to tune the simulation parameters. To keep the example simple, we compared real and simulated data using four metrics: average distance per trip (ADT), average distance per agent per day (ADA), maximum trip distance (MXD), and median trip distance (MDD). The similarity score in Equation 1 quantifies the match, with values closer to 1 indicating greater similarity to GeoLife.
@@
To calibrate simulations, we developed a genetic algorithm integrated into the data generation framework. We identified 63 parameters affecting agent behavior (e.g., number of interests, walking speed, rental cost ratio). A full list and documentation are provided in our GitHub repository. The goal was to find parameter values producing trajectories that best match GeoLife metrics. The algorithm begins with simulations of $n$ using valid parameter values randomly chosen. Each run is scored using Equation 1, and the best performing parameter sets are retained as “parents.” New parameter sets (“children”) are then generated by combining parent values in five ways: maximum, minimum, mean, random mix, or random mutation. This process repeats until $n$ children are produced, each followed by a simulation and scoring step. New generations are iteratively created until convergence or manual termination. The best performing parameters across all generations are selected as the calibration result. The implementation is publicly available on GitHub.
@@
The raw outputs generated by the simulator must be transformed into structured, application ready datasets through a modular and scalable processing pipeline. After data generation, simulation logs can be processed in two primary ways: (1) by concatenating the complete set of split log files, or (2) by trimming and merging logs from a selected subset of simulation instances. In both cases, standardization is often required. For example, spatial coordinates recorded in the agentStateTable are not stored in standard GPS format and must be converted before downstream use. Our processing scripts convert these raw logs into structured datasets suitable for tasks such as machine learning model training, calibration, and statistical analysis. The agentStateTable, which records the state of each agent at every simulation tick, can be split into manageable chunks, filtered by relevant fields, and reassembled into a unified dataset. Users can explicitly specify which fields are retained, enabling targeted data extraction tailored to specific research objectives.
@@
HD-GEN is designed with scalability and customizability as core principles. To reduce unnecessary log generation, we provide detailed documentation that guides users in configuring the simulator to emit only essential fields such as timestamps, geographic coordinates, and agent identifiers. This design minimizes storage and computational overhead while improving overall efficiency. For very large datasets, the pipeline includes Python scripts capable of processing hundreds of gigabytes of data, as well as lightweight Bash utilities that perform file system level concatenation and column extraction. In addition to filtering redundant fields, the pipeline supports embedding ground truth labels and other annotations, ensuring that generated datasets are well aligned with downstream research and application requirements.
@@
Another critical aspect of HD-GEN is the visualization of processed data. Visualization offers a powerful means of interpreting simulation outputs by uncovering patterns, anomalies, and behavioral dynamics that are often difficult to detect in raw numerical logs. It is particularly valuable for validating the effects of parameter adjustments, identifying unexpected or emergent agent behaviors, and improving the overall accuracy of the model. When new features or anomalous behaviors arise, visual inspection provides an intuitive method for confirming their presence and assessing their impact across both temporal and spatial dimensions.

### v2

@@ Abstract.
Understanding individual-level human mobility is critical for a wide range of applications. As such, real-world trajectory datasets provide valuable insights into actual movement behaviors and patterns of life but are often constrained by data sparsity and participant bias. Synthetic data, by contrast, offers scalability and flexibility but frequently lacks realism. To address this gap, we introduce a comprehensive software pipeline for, generating, calibrating, processing, and visualizing large-scale individual-level human mobility datasets that combine the realism of empirical data with the control and extensibility of Patterns-of-Life simulations. Our system consists of four integrated components. (1) a data generation engine which constructs geographically grounded simulations using OpenStreetMap data to produce diverse mobility logs. (2) a genetic algorithm–based calibration module that fine-tunes simulation parameters to align with real-world mobility characteristics, such as daily trip counts and radius of gyration, enabling realistic behavioral modeling. (3) a data processing suite which transforms raw simulation logs into structured formats suitable for downstream applications, including model training and benchmarking, and (4) a visualization module that extracts key mobility patterns and insights from the processed datasets and presents them through intuitive visual analytics for improved interpretability.
Trajectory Generation, Patterns of Life, Simulation, Trajectory Data
@@ 1. Introduction
Human mobility datasets are fundamental for understanding human behavior (Toch et al., 2019; Zhu et al., 2024a; Gonzalez et al., 2008), mobility patterns (Mokbel et al., 2024; Schneider et al., 2013), and traffic dynamics (Chen et al., 2022; Islam et al., 2021). They enable a wide range of applications, including outlier and anomaly detection (Zhang et al., 2024a, b; Amiri et al., 2024b), infectious disease modeling (Kohn et al., 2023), and urban planning and infrastructure design (Isaacman et al., 2012). Mobility datasets can be broadly categorized as real or synthetic. Real world data, such as those collected from cell phone traces, provide high fidelity but raise serious quality and privacy concerns, while synthetic data offer improved scalability and privacy preservation at the cost of reduced realism (Züfle et al., 2024). In practice, privacy regulations (Xiao and Xiong, 2015; Khoshgozaran et al., 2011; Krumm, 2009), de-identification techniques, and voluntary data collection introduce sparsity and information loss, resulting in significant trade-offs between data utility, realism, and accessibility (Zhu et al., 2024b; Amichi et al., 2025).
@@
To mitigate these limitations, researchers are increasingly relying on simulated trajectory datasets which provide scalable, noise free mobility information without privacy risks. However, many of the existing trajectory simulators rely on random or simplistic destination selection (Düntgen et al., 2009; Brinkhoff, 2002) or parametric trip distributions (Levandoski et al., 2012; Armenatzoglou et al., 2013). While such approaches are suitable for benchmarking storage systems or query processing, we would argue that they fail to capture the intentional and routine driven nature of real human mobility.
@@
The Pattern of Life (POL) simulation framework (Kim et al., 2020; Züfle et al., 2023; Amiri et al., 2024a) addresses this gap by modeling human needs, activities, and daily routines. POL generates purposeful mobility behaviors such as commuting, dining, and social interactions, producing trajectories that better reflect real world movement semantics (e.g., going to and from work). POL has been used to generate large scale datasets with realistic mobility patterns, social interactions, and check-in behavior. Prior work includes the generation of fine grained mobility and social network datasets across multiple U.S. cities (Amiri et al., 2023), datasets with injected anomalies for mobility anomaly detection (Amiri et al., 2024b), and a GeoLife calibrated dataset optimized via genetic algorithms to balance realism and scalability (Amiri et al., 2024c).
@@
In this work, we extend our SIGSPATIAL 2025 short paper (Amiri et al., 2025) and consolidate our prior contributions (Amiri et al., 2024a, c, 2023) into a complete and systematic POL based data generation system. We introduce a unified software platform that integrates trajectory generation, calibration, processing, and visualization into a coherent end to end pipeline, referred to as HD-GEN (i,e., Human mobility Data GENeration based on POL) . The platform expands the functionality of POL while emphasizing reproducibility, scalability, and usability for large scale experimental and operational settings.
@@
  • delivers an improved open source repository ({https://github.com/onspatial/hd-gen}) with clearer documentation and integrated visualization toolkits.
The remainder of the paper is organized as follows. In Section 2, we review relevant background information and related prior work. Next , in Section 3, we provide a technical overview of POL and describe the software design and system-level details of HD-GEN. In Section 4, we summarize the datasets produced using HD-GEN. Finally, in Section 5, we conclude by summarizing the contributions of this work and discussing directions for future research.
2. Background Information and Related Work
In this section, we provide background information and related work on human mobility data generation, including an overview of the POL simulation and a discussion of existing trajectory datasets. We first describe the POL framework, its behavioral foundations, and its suitability for generating large-scale, realistic mobility data. We then briefly review commonly used real and synthetic trajectory datasets and highlight their limitations, motivating the use of POL and the need for calibration against observed mobility patterns.
@@ 2.1. The Patterns-of-Life Simulation (POL)
The POL simulation, introduced in Urban Life (Züfle et al., 2023), is a city-scale, agent-based modeling framework designed to generate realistic human mobility trajectories and interaction patterns. The simulation instantiates autonomous agents within an urban environment constructed from OpenStreetMap data, including road networks, residential zones, workplaces, and recreational venues. Agents perceive and interact with this environment and each other while independently making decisions that reflect the routines and choices of individuals in everyday life.
Agent behavior in POL is governed by a combination of physiological, financial, and social needs inspired by Maslow’s hierarchy of needs (Maslow, 1943), together with goal-oriented decision making based on the Theory of Planned Behavior (Ajzen, 1991). At each simulation time step, agents evaluate competing needs such as earning income, acquiring food, maintaining social relationships, and participating in leisure activities. These needs are balanced against practical constraints including time availability, spatial accessibility, and personal resources. The resulting decisions determine when and where agents travel, producing realistic daily schedules and mobility patterns. As agents move between different types of locations, the simulation generates fine-grained spatiotemporal trajectories as well as implicit social graphs that emerge from repeated co-location and interaction. This design makes POL well suited for controlled experimentation and downstream analysis in mobility modeling, social network analysis, and urban studies.
@@
Fig. 1 illustrates the interaction between a person agent and the simulation world. Each person is modeled as an autonomous agent with multiple needs, including food, sleep, shelter, financial safety, and love need from social connections. These needs evolve continuously over time and drive the agent’s decision making process. To satisfy its needs, the agent interacts with the environment, which defines mobility constraints, a spatial network, and points of interest such as workplaces, homes, and public venues. The environment determines where the agent can move and which actions are feasible at each location. Each agent perceives their environment, selects actions based on its current needs and context, and executes those actions through movement and interactions with points of interest, such as working at a workplace, or with other agents at recreational sites. All agent states, actions, and environment responses are recorded by the log system. The generated logs capture realistic spatiotemporal behavior and serve as the output of the simulation for downstream analysis and modeling.
@@
The POL simulation is implemented in Java utilizing the MASON toolkit (Luke et al., 2018) and is configurable through a command-line interface, and publicly available on GitHub (Züfle et al., 2023). Users can specify city regions, population sizes, and behavioral parameters, enabling reproducible and extensible experiments. Fig. 2 presents the end to end simulation workflow. The process begins with a parameter sweep, where multiple configurations are defined to explore variations in agent behavior and environmental conditions. These configurations are formalized during parameterization, producing a consistent set of parameters that govern the simulation. In parallel, the map generator produces spatial representations of the environment, including networks and points of interest. The generated maps, together with the selected parameters, are passed to the initialization stage, where agents, environments, and initial states are instantiated. The simulation stage executes agent interactions over time based on the initialized state and active parameters. During execution, manipulation modules can dynamically modify parameters or system states, enabling interventions, scenario changes, or policy testing. The simulation state can be serialized to disk to implement checkpointing, allowing execution to be paused, restored, and reproduced in long running experiments. All simulation outputs are stored in structured formats guided by a predefined schema definition, which specifies how agent states, actions, and environmental events are recorded. The log system collects these records, producing standardized logs that serve as the final output of the simulation.
@@ 2.2. Related Work and Relation to HD-GEN
Human mobility data can be collected from real-world observations or synthesized using computational models. Real-world trajectory datasets capture observed mobility behavior and provide valuable insight into human movement and activity patterns. A widely used example is the GeoLife dataset (Zheng et al., 2011), which contains four years of high-resolution GPS traces from approximately 180 users in Beijing, spanning both commuting and leisure activities. More recently, the large-scale YJMOB100K dataset (Yabe et al., 2024) provides anonymized mobility records from 100,000 users at 30-minute intervals, aggregated into 500-meter spatial cells. While this scale enables population-level analysis, the spatial aggregation limits fine-grained, place-based behavioral inference. Additional trajectory datasets focus on vehicles, including taxis (Yuan et al., 2010; Piorkowski et al., 2009) and buses (Dias and Costa, 2018). These datasets are valuable for modeling transportation systems but do not accurately represent individual human activity patterns, since each trip may involve different occupants and does not reflect continuous personal behavior. Synthetic trajectory datasets address limitations of real mobility data, including privacy constraints, proprietary access, and restricted scale. Recent deep generative approaches learn mobility patterns to produce realistic spatiotemporal trajectories, including factorized models such as EETG (Zhang et al., 2022) and diffusion-based methods (Zhu et al., 2024b). Optimization-based techniques further aim to enhance incomplete or sparse trajectory data (Zhu et al., 2024a; Hu et al., 2022). In contrast, the POL simulation framework (Kim et al., 2020; Züfle et al., 2023; Amiri et al., 2024a) generates trajectories driven by modeled human needs and decision processes rather than purely geometric or kinematic realism.
In our prior work (Amiri et al., 2023), POL was used to generate a large-scale dataset exceeding 1.5 TB of mobility, check-in, and social interaction data. We further demonstrated the resulting data generator and its ability to reproduce key mobility and activity patterns in (Amiri et al., 2024a). Although the default parameter settings yield qualitatively realistic behavior, regional variation was not systematically characterized. This limitation motivated our subsequent study (Amiri et al., 2024c), in which we calibrated POL to better match mobility patterns observed in the GeoLife dataset for Beijing. Finally, we integrated these efforts into a unified framework, HD-GEN, which was introduced in a concise system paper (Amiri et al., 2025).
@@ 3.2. POL Details
This section describes the new features added to POL as part of this papers extension, and the changes made to the original POL simulation to improve its performance. Additionally, we explain the checkpointing, which allows for saving and resuming simulation states, and the log streaming feature that enables real-time monitoring of simulation progress.
3.3. Performance Optimization
Table 1. Performance comparison between Vanilla and Enhanced simulations for a 10 simulation day run. Initialization and simulation times are reported for different numbers of agents. A value of -1 indicates that the simulation did not complete. (Init: Initialization time; Sim: Simulation time; Exit On/Off: agent can/cannot exit; s: seconds; m: minutes; h: hours)
Agents Vanilla Exit On Vanilla Exit Off Enhanced Exit On Enhanced Exit Off
Init Sim Init Sim Init Sim Init Sim
1000 33.2 s 10.3 m 30.7 s 15.2 m 8.7 s 1.3 m 4.9 s 1.0 m
2000 1.9 m 36.2 m 1.4 m 51.5 m 11.6 s 2.0 m 11.1 s 2.0 m
3000 2.6 m 1.4 h 2.2 m 1.9 h 19.1 s 3.1 m 20.9 s 3.1 m
4000 4.2 m 3.0 h 4.1 m 3.6 h 35.3 s 4.8 m 35.6 s 5.3 m
5000 4.6 m 3.7 h 5.7 m 4.3 h 48.8 s 5.7 m 53.0 s 5.8 m
10000 13.9 m 15.8 h 13.0 m 16.7 h 3.3 m 14.7 m 3.3 m 14.7 m
15000 24.1 m 35.4 h 22.8 m 41.2 h 8.7 m 32.6 m 8.0 m 28.9 m
20000 -1 -1 -1 -1 12.8 m 40.7 m 11.2 m 36.9 m
25000 -1 -1 -1 -1 18.9 m 56.7 m 19.3 m 58.0 m
50000 -1 -1 -1 -1 1.1 h 2.7 h 1.0 h 2.7 h
100000 -1 -1 -1 -1 4.4 h 9.0 h 4.5 h 9.4 h
150000 -1 -1 -1 -1 9.6 h 19.3 h 9.8 h 20.1 h
As we aim to simulate larger populations within the Pattern of Life (POL) simulation, performance optimization becomes a critical concern. The original Vanilla POL simulation exhibited limitations in scalability, particularly as the number of agents increased. To address these challenges, we implemented a series of performance optimizations in the Enhanced POL simulation. The Enhanced simulation incorporates several key optimizations, including optimize the initial parameter values, avoid logging redundant or unnecessary information, and replace the optimal travel distance algorithm with a Euclidean distance calculation (Khan et al., 2024) to identify the nearest location. These changes collectively contribute to a more efficient simulation process, reducing both initialization and execution times.
To evaluate the effectiveness of the proposed performance optimizations, we focus on the execution time of a single simulation instance. Since parallel execution does not alter the runtime of an individual simulation, this evaluation isolates the impact of the internal optimizations introduced in the enhanced system. All experiments were conducted by simulating a period of 10 days (14,400 simulation ticks) while varying the number of agents. The experiments were executed on a machine equipped with an Intel Core i7-6700HQ CPU and 24 GB of RAM. We compare the baseline Vanilla simulation against the Enhanced simulation, which incorporates the proposed performance optimizations and we will use the Enhanced simulation as the default simulation in HD-GEN going forward unless otherwise specified.
For each configuration, we report two distinct performance phases:
  • Initialization time, which measures the duration required to load spatial data, instantiate agents, initialize environments, and prepare simulation state.
  • Simulation time, which measures the wall-clock time required to execute the full 10-day simulation period.
Results are reported for four execution modes: Vanilla and Enhanced simulations, each evaluated with agent exit conditions enabled (exit_on) and disabled (exit_off). The exit condition allows agents to leave the simulation when predefined criteria are met, reducing computational load in some scenarios. A value of -1 indicates that the simulation did not complete within a reasonable time frame or was manually terminated due to excessive runtime.
Table 1 summarizes the results across a wide range of agent counts. The results demonstrate that the Vanilla simulation fails to scale beyond approximately 15,000 agents, frequently exceeding practical execution limits. In contrast, the Enhanced simulation completes successfully for populations exceeding 150,000 agents, while substantially reducing both initialization and simulation times across all tested configurations. These results confirm that the proposed optimizations significantly improve scalability while preserving the behavioral semantics of the underlying simulation.
@@ 3.3. Performance Optimization
The POL simulation is deterministic, meaning that executing the simulation with the same input parameters always produces identical results. The simulation is also portable across different machines and operating systems, allowing execution to continue without breaking the pipeline or altering outcomes. These properties are essential for exploring multiple scenarios while preserving previously executed configurations.
@@ 3.4. HD-GEN Details
To generate simulated data, we developed tools for regional map preparation, simulation execution, and large scale dataset generation in parallel. Map generation relies on OpenStreetMap (OSM) data accessed through the Overpass API (API, 2024) to extract building footprints, road networks, and land use information. These raw features are further refined using predictive models to classify building usage. The map generation process produces three shapefiles, buildings.shp, buildingUnits.shp, and walkways.shp, which collectively define the spatial environment used by the POL simulation. Figure 6 shows an example script that generates a new map for the Minneapolis region. In addition to bounding box based queries, the tools support custom Overpass queries, allowing flexible selection of geographic regions and features for map generation.
@@
For calibration, we used the GeoLife (Zheng et al., 2010, 2011) dataset as a real-world reference and applied a genetic algorithm guided by a similarity function to tune the simulation parameters. To keep the example simple, we compared real and simulated data using four metrics: average distance per trip (ADT), average distance per agent per day (ADA), maximum trip distance (MXD), and median trip distance (MDD). The similarity score in Equation 1 quantifies the match, with values closer to 1 indicating greater similarity to the GeoLife dataset.
@@
To calibrate the simulations, we developed a genetic algorithm integrated into the data generation framework. We identified 63 parameters affecting agent behavior (e.g., number of interests, walking speed, rental cost ratio). A full list and documentation are provided in our GitHub repository ({https://github.com/onspatial/hd-gen}). The goal was to find parameter values producing trajectories that best match GeoLife metrics. The algorithm begins with simulations of $n$ using valid parameter values randomly chosen. Each run is scored using Equation 1, and the best performing parameter sets are retained as “parents.” New parameter sets (“children”) are then generated by combining parent values in five ways: maximum, minimum, mean, random mix, or random mutation. This process repeats until $n$ children are produced, each followed by a simulation and scoring step. New generations are iteratively created until convergence or manual termination. The best performing parameters across all generations are selected as the calibration result. The implementation is publicly available on GitHub.
@@
The raw outputs generated by the simulator need to be transformed into structured, application ready datasets which is done through a modular and scalable processing pipeline. After data generation, simulation logs are processed in two primary ways: (1) by concatenating the complete set of split log files, or (2) by trimming and merging logs from a selected subset of simulation instances. In both cases, standardization is often required. For example, spatial coordinates recorded in the agentStateTable are not stored in standard GPS format and must be converted before downstream use. Our processing scripts convert the raw logs into structured datasets suitable for tasks such as machine learning model training, calibration, and statistical analysis. The agentStateTable, which records the state of each agent at every simulation tick, can be split into manageable chunks, filtered by relevant fields, and reassembled into a unified dataset. Users can explicitly specify which fields are retained, enabling targeted data extraction tailored to specific research objectives.
@@
HD-GEN is designed with scalability and customizability as core principles. To reduce unnecessary log generation, we provide detailed documentation on our Github page that guides users in configuring the simulator to emit only essential fields such as timestamps, geographic coordinates, and agent identifiers. This design minimizes storage and computational overhead while improving overall efficiency. For very large datasets, the pipeline includes Python scripts capable of processing hundreds of gigabytes of data, as well as lightweight Bash utilities that perform file system level concatenation and column extraction. In addition to filtering redundant fields, the pipeline supports embedding ground truth labels and other annotations, ensuring that generated datasets are well aligned with downstream research and application requirements.
@@
Another critical aspect of HD-GEN is the visualization of processed data. Visualization offers a powerful means of interpreting simulation outputs by uncovering patterns, anomalies, and behavioral dynamics that are often difficult to detect in raw numerical logs (Grimm, 2002). It is particularly valuable for validating the effects of parameter adjustments, identifying unexpected or emergent agent behaviors, and improving the overall accuracy of the model. When new features or anomalous behaviors arise, visual inspection provides an intuitive method for confirming their presence and assessing their impact across both temporal and spatial dimensions.